OpenStudy (anonymous):

Find the sixth term of a geometric sequence with t5 = 24 and t8 = 3

OpenStudy (anonymous):

Usually, ration would be given, so you could plug it into the formula: a(n) = a(1)r^(n-1)

OpenStudy (anonymous):

*ratio

OpenStudy (anonymous):

does the sign mean ratio?

OpenStudy (anonymous):

t5 = a*r^4 = 24 t8 = a.r^7 = 3 Dividing one by the other r^3 = 1/8 r = 1/2 t6 = a.r^5 = t5*r = 24 * 1/2 = 12

OpenStudy (anonymous):

The symbol to the right of the 'a' is a multiplication symbol.

OpenStudy (anonymous):

oh okayy couldnt tell.. thank you!

OpenStudy (anonymous):

No prob :)